the length of a rectangle is 2 inches greater than twice the width and it has an area of 24 square feet, find its dimensions ?
Answer by stanbon(75887)

the length of a rectangle is 2 inches greater than twice the width and it has an area of 24 square feet, find its dimensions ?
------
Let the width be "x"::
Then length is "2x+2"
=================
Equation: Area = length*width
24 = x(2x+2)
12 = x(x+1)
x^2 + x - 12 = 0
Factor:
(x+4)(x-3) = 0
Positive solution:
x = 3 inches (width)
2x+2 = 8 inches (length)
=========================
Cheers,
Stan H.
-------------------------